OpenStudy (anonymous):

How do I do this and what is the answer? Write the equation of the line that is parallel to the line 4x - 3y = -12 and passes through the point (-3, 4). A) y = 4/3 x + 8 B) y = 4/3 x + 3 C) y = - 3/4 x +8 D) y = - 3/4 x + 3

OpenStudy (anonymous):

Well you would put the equation into slope intercept form first.

OpenStudy (anonymous):

Do you know how to do that? @jut21

OpenStudy (anonymous):

no

OpenStudy (anonymous):

@teagirl7630

OpenStudy (anonymous):

ah okay. So to get the equation into slope intercept form, you have to get the y alone. The equation for slope intercept form is y=mx+b. Your original equation is 4x-3y=-12. The way to get this into slope intercept form is to isolate the y. Do you understand how to do that? @jut21

OpenStudy (anonymous):

not really... sorry

OpenStudy (anonymous):

That's okay! So the original equation is: 4x-3y=-12. So the first thing when isolating the y is to move the 4x to the other side. So the equation then looks like this: -3y=-4x-12. You want y alone, so you have to divide both sides of the equation by -3. The equation would then look like this : y=4/3x+4. Does that make sense?

OpenStudy (anonymous):

@jut21

OpenStudy (anonymous):

So then the answer to the question would be B) y=4/3x+3. Oh and the slope intercept equation is y=mx+b. And the answer to the question satisfies that. :) If you need more explanation just tag me in your question again.

OpenStudy (anonymous):

@teagirl7630 are you sure that the answer is not A)?